OpenStudy (anonymous):

suppose lim[x:4,f(x)=4] and lim[x:4,g(x)= -2], find, lim[x:4,f(x)]

hartnn (hartnn):

@adosa Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) are you sure you need, lim[x:4,f(x)] ?? because as per your question, you are given lim[x:4,f(x)]=4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

suppose lim[x:4,f(x)=4] and lim[x:4,g(x)= -2], find, lim[x:4,xf(x)]

hartnn (hartnn):

okk, so you need \(\lim \limits_{x \rightarrow 4}x f(x)\)

OpenStudy (anonymous):

yes

hartnn (hartnn):

from \(\lim \limits_ {x\rightarrow 4}f(x)=4 , \\ \implies f(4)=4\)

hartnn (hartnn):

so, \(\lim \limits_{x \rightarrow 4}x f(x)= 4 f(4)=.... ?\)

OpenStudy (anonymous):

suppose lim[x:4,f(x)=4] and lim[x:4,g(x)= -2], find, lim[x:4,g(x)/f(x)-1]

hartnn (hartnn):

did you get the answer to lim[x:4,xf(x)] ??

OpenStudy (anonymous):

4 right?

hartnn (hartnn):

no... 4 f(4) = 4*4 =16

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

suppose lim[x:4,f(x)=4] and lim[x:4,g(x)= -2], find, lim[x:4,g(x)/f(x)-1]